# Gas Cost Analysis of Proxy and Diamond Patterns: Towards Trusted Smart Contract Engineering in EVM Blockchains

### v1

@@ 4.3 Results Retrieval
In addition, the framework furnishes complete stack traces of the calls, detailing the gas consumption for each operation. These traces are utilized to construct charts that capture the gas cost of every individual function call. This level of detail supplements gas reports, which typically provide only a summary of costs, including the minimum, average, median, and maximum values.
@@ 5 Evaluation
This section presents the gas costs evaluation of the Proxy and Diamond smart contract patterns, against a baseline built using the classic pattern in the context of an app deployment and upgrade  [^5] All results can be found here: https://github.com/opixelum/proxy-diamond-patterns-gas-analysis/tree/main/data.
5.1 Gas Cost During Upgradeable Smart Contract Deployment and Upgrades
Figure 3: Deployment Cost by Pattern & Version.
@@ 5.1 Gas Cost During Upgradeable Smart Contract Deployment and Upgrades
We first aim at comparing the gas costs during the deployment (Version 1) and upgrades (Version 2 and 3) of the file notarization application implemented with the three different patterns. For the Classic and Proxy pattern, a new contract is deployed for each version. For the Proxy, the proxy pointer is updated in the smart contract implementation. For the Diamond pattern, in Version 2, only facet is changed and Version 3, all facets are changed. Figure 3 presents the deployment costs, in gas units, of each version by pattern. The cost of deployment increases with each version, except for the Diamond pattern, where the initial cost of deployment is significant. In total, the classic pattern emerges requires 1,614,545 gas units, while the Proxy and Diamond patterns consume approximately 2.6 times more, with 4,343,104 and 4,123,977 gas units respectively. The classic pattern appears the most gas efficient. The increased consumption at deployment of the Proxy and Diamond patterns relates to the need to deploy more contracts compared to the classic pattern: two for the Proxy pattern, and four for our example with the Diamond pattern.
5.2 Gas cost during upgradeable smart contract execution: study of all functions
Figure 4: Average Function Cost by Pattern for Version 2
The analysis now turns to comparing the average gas cost of each function, for seeing which pattern is the most gas efficient when it comes to execution. Figure 4 depicts the gas costs observed for Version 2 averaged per function. It is to note that we observe the same gas costs behavior in Version 1 and 3 (c.f., online gas reports). The addFile function is the costliest, followed by the updateFile function. The other functions require less gas. There is no significant difference in gas cost among the three patterns. The Proxy pattern is slightly more expensive than the classic pattern, and the Diamond pattern is slightly more expensive than the Proxy pattern. The most important difference is for compareHashes being 300 times cheaper than addFile for the classic pattern.
The difference in gas costs between compareHashes and addFile is due to the type of operations done in the functions. The reason is that the addFile does two writings in the storage, whereas the updateFile does only one. The other functions only do reads in the storage, which cost less gas than writes.
The classic pattern does not delegate any call to another contract, whereas the Proxy and Diamond patterns do, so it adds a cost. Then, when calling a diamond, it needs to do more operations than a proxy to find the right facet to call. This issue is not present with the Proxy pattern, as there is only one implementation contract. Now, the differences between the view and pure functions are examined. getFile... functions are view functions, which means that they only do a state lookup. This lookup can cost up to 2100 gas for a cold storage load (when reading a variable for the first time) [1]. compareHashes is a pure function, which means that it does not do any state lookup or modification. It only does some computations on given parameters, blockchain’s global variables or not. This is why it requires less gas than getFile... functions in the classic pattern.
5.3 Gas cost during upgradeable smart contract execution: more in-depth study of the addFile function
Figure 5: Gas Cost of the addFile Function by File Name Configuration
To gain a deeper understanding of the gas cost implications, the addFile function is subjected to three specific test cases, each with different configurations for the file name parameter to provide a more detailed analysis of how various input scenarios impact gas consumption. Figure 5 provides the results of the experiment. For the classic pattern, gas cost increases after each 32 characters, and remains the same for each iteration. Gas cost is expensive for the first iteration, then has an important drop for all the other iterations (it stays the same for each remaining iteration). Results are similar for the Proxy and Diamond patterns. The results for the first configuration can be explained bt the storage slot that is 32 bytes long. When a storage slot is full, the EVM needs to use another storage slot, which has a cost. That is why in the second configuration, the gas cost is the same for each iteration, since the length of the file name is the same. But because the last character is different, the EVM needs to write in the storage, which has a cost. Finally, the last configuration is way cheaper than the previous one after the first iteration because the EVM does not need to edit the storage, since the file name is the same.
5.4 Summary of Comparative Test Findings
@@
In reviewing the data obtained from the comparative analyses, it is evident that the deployment phase exhibits the most significant variation among the different patterns, primarily due to the fluctuating number of smart contracts deployed. During execution, the gas costs are relatively similar across the patterns, though there is a marginal escalation in costs when transitioning from the classic architecture to the Proxy pattern, and subsequently from the Proxy to the Diamond pattern. However, this incremental increase in gas costs may be considered negligible, depending on the specific requirements of the software.
6 Theory Generation
6.1 Proxy and Diamond Patterns
We propose to summarize and consolidate the Proxy and Diamond blockchain patterns by following the Alexandrian form format proposed by Christopher Alexander [3].
6.1.1 Proxy Pattern
@@
  • Summary - The Proxy Pattern facilitates upgradability in smart contracts through a Proxy contract pointing to the latest version of a logic contract.
@@
  • Context - A smart contract must be upgraded due to evolving requirements and potential improvements [9].
@@
  • Problem - Traditional smart contracts lack the ability to be updated without manual storage migration, posing challenges in addressing vulnerabilities, enhancing functionality, and adapting to changing circumstances.
  • Forces
    • Immutability vs. Upgradability - Smart contracts on blockchain platforms are traditionally immutable once deployed, ensuring security but hindering adaptability.
    • Gas Costs vs. Flexibility - The balance between minimizing gas costs and providing flexibility for contract upgrades is a critical challenge.
    • Trust vs. Transparency - Establishing trust in the upgrade process while maintaining transparency is a delicate balance.
@@
  • Solution - The Smart Contract Proxy Pattern introduces a proxy contract as an intermediary layer. This proxy delegates calls to a logic contract, allowing seamless upgrades by deploying new logic contracts and updating the proxy to point to the latest version, hence allowing for dynamic updates.
  • Consequences
@@
        Upgradability: Enables systematic and non-disruptive upgrades by employing a distinct proxy contract that separates logic and state. Manual data migration is not required, as the storage is maintained within the proxy contract. This allows the logic contract to be updated without altering the state-holding proxy contract.
        Seamless Interaction with Existing State: Maintains the contract’s address and state during upgrades when the logic contract in the Proxy pattern is upgraded. The new version can seamlessly interact with the existing state, minimizing the need for explicit data migration.
        Ease of implementation: Several libraries offer an easy integration of this pattern into smart contracts, encapsulating the proxy management (e.g., Open Zepplin and EIP-897 implementations)
@@
        Proxy Management Functions Complexity: Requires dedicated proxy management functions, contingent on the type of proxy used (transparent or UUPS). This introduces an extra layer of complexity in the development process, demanding specific knowledge and consideration during implementation.
        Compatibility Maintenance: Ensuring compatibility between different contract versions in the Proxy pattern requires meticulous consideration. Preserving consistent function selectors and storage layouts becomes crucial for a smooth transition during upgrades.
        Limited Direct Function Visibility: Indirect access to the logic functions visibility. Users only have a direct access to the proxy administration functions, necessitating consultation of accompanying documentation to identify callable functions. This may potentially introduce usability challenges for end-users.
  • Related Patterns - Diamond Pattern and upgradeable pattern variations [8, 4]
  • Known Uses
    • EIP-897: ERC-Proxy Standard: An Ethereum Improvement Proposal outlining a standardized implementation of proxy contracts[23].
    • OpenZeppelin’s Proxy Patterns: Practical implementations and documentation in popular smart contract development frameworks[4].
    • Compound and USDC: Respectively a DeFi protocol and a stable coin leveraging the Proxy Pattern for upgrades
6.1.2 Diamond Pattern
  • Summary - The Diamond Pattern addresses the challenge of contract size limitations and enhances maintainability and versioning by employing multiple implementation contracts.
  • Context - Need for a solution to improve maintainability in large contracts.
  • Problem - Traditional approaches face challenges in managing contract size and versioning effectively, limiting the scalability and maintenance of smart contracts. The need for more scalable and maintainable smart contract structures is evident.
  • Forces
    • Contract Size vs. Modularity - Balancing the need for compact contract sizes with the demand for modular, well-organized code structures poses a challenge.
    • Maintainability vs. Simplicity - Achieving improved maintainability without introducing unnecessary complexity is an ongoing struggle.
    • Scalability vs. Consistency - Scaling smart contract applications can hinder consistency as it requires accommodating versioning and updates.
  • Solution - The Diamond Pattern introduces a structure where a proxy can point to multiple logic contracts. The Diamond pattern deployment involves the deployment of all contracts (diamond and facets), retrieving facets function selectors, and leveraging both to implement a diamond cut.
  • Consequences
@@ 6.1.2 Diamond Pattern
        Structured and Modular design: Allows for multiple logic contracts (facets) which facilitate large contracts code organization.
        Flexible maintenance: Easy addition, replacement, or removal of facets, facilitating smoother upgrades and interactions with existing data, as well as scalable versioning.
        Compilation time for facet updates: The modular facet-based code organization shortens compilation time, as editing a facet only requires the compilation of that specific facet.
@@
        Implementation Complexity: Complexity in smart contract variables storage management and facet-library interactions. This increased complexity demands specific knowledge and careful consideration during the implementation phase as there is currently a lack of supporting implementation libraries.
        Complexity in managing multiple logic contracts: The Diamond pattern’s enhanced flexibility comes at the cost of increased complexity, necessitating meticulous planning and testing to ensure smooth data interaction and integrity across different contract versions. The intricacies of managing multiple logic contracts require careful consideration to avoid unintended consequences during upgrades.
  • Related Patterns - Proxy Pattern and upgradeable pattern variations [8, 4]
  • Known Uses
    • Ethereum Improvement Proposal EIP-2535 [8]
    • Aavegotchi: a Non-Fungible Token (NFT) based gaming protocol
    • GeoWeb: a dApp managing digital land property rights using NFTs
6.2 Choosing between Proxy and Diamond Patterns
Figure 6: Decision model for Upgradeable Smart Contract Pattern Usage
Figure 6 proposes a decision model to orient between a Proxy or a Diamond pattern in the design stage of an upgradeable smart contract. It follows the design goal decision model introduced by Xu  [32] where decisions are modeled using a logical BPMN flow. More precisely, arrows, logical gateways, and functional and non-functional properties orient the decision path.
The choice of smart contract pattern largely depends on the need for upgradability. For scenarios without upgradability needs, the classic pattern is preferable due to its straightforward development process, relying on contract inheritance and library imports. However, upgrades in the classic pattern often involve complex and resource-intensive data migration. This pattern also poses challenges in communicating new contract addresses to users, potentially affecting user experience. For extensive upgradeable features, the Diamond Pattern is recommended. Its modular nature allows for easy addition or removal of facets, reducing compilation time. However, it is less cost-effective initially compared to the Proxy pattern and requires in-depth knowledge of smart contract storage and facet-library management. The Proxy pattern is advised for limited code sizes or infrequent upgrades. It simplifies development and integrates easily with libraries like OpenZeppelin’s. This pattern enhances upgradability by separating logic and state, reducing the need for data migration. But, it offers less flexibility and modularity compared to the Diamond Pattern and demands careful consideration to maintain compatibility across versions.
In conclusion, while the classic pattern excels in execution ease, the Proxy and Diamond patterns provide a more manageable framework for upgrades, simplifying contract interactions for users during updates.
7 Discussion and Conclusion
@@
Each pattern unfolds with distinct strengths and weaknesses, delineating its applicability across diverse scenarios. The Diamond pattern exhibits modularity countered by higher initial costs, and the Proxy pattern strikes a balance between the classical approach (without proxy) and the Diamond one.
To generalize these initial findings, the study advocates for extending experiments beyond the notarization scenario used as a comparison baseline for this paper. Essential to this endeavor is the necessity for additional experiments encompassing diverse blockchain networks and pattern libraries to extrapolate and validate the findings. Moreover, a replication of the study on alternative upgradeable patterns would also provide a more comprehensive understanding of upgradeable smart contract patterns.
Despite these considerations, real-world implementations in projects like Compound, USDC, GeoWeb, and Aavegotchi underscore the substantial impact of the Proxy and Diamond patterns in the blockchain domain. These patterns serve as facilitators of flexible, upgradeable, and scalable smart contracts, showcasing their adaptability across a spectrum of blockchain applications. In the dynamic landscape of the blockchain industry, the strategic selection and adept implementation of these design patterns emerge as pivotal determinants of success and innovation in decentralized applications.
Looking ahead, future research should delve into the dynamic nature of these patterns, potentially exploring hybrid upgradability models or introducing novel paradigms in smart contract development. Additionally, a deeper exploration into blockchain sharding and layer 2 solutions holds promise for offering valuable insights into the scalability of these patterns in diverse blockchain ecosystems.

### v2

@@ 4.4 Results Retrieval
These gas cost results for each function are derived by summing the gas costs associated with every operation performed within the function. Each operation, also known as opcode, has a fixed gas cost, which is the same for all patterns. Those instructions are described in the Ethereum Yellow Paper [31]. For example, the SLOAD opcode, which reads a value from storage, has a fixed cost of 100 for warm access, and 2100 gas for cold access. Then, the deployment cost is the sum of several components. First, the TRANSACTION opcode incurs 21,000 gas units, representing the base cost of every transaction on the EVM. Additionally, there is the CREATE opcode, costing 32,000 gas units, which is used for creating a new contract. Next, the cost related to the bytecode includes 4 times the number of 0 bytes and 16 times the number of non-zero bytes. Furthermore, 200 gas units are added for every byte of the contract’s size. Finally, if a constructor function is present, its cost is also included in the deployment cost.
In addition, the framework furnishes complete stack traces of the calls, detailing the gas consumption for each operation. These traces are utilized to construct charts that capture the gas cost of each function call. This level of detail supplements gas reports, which typically provide only a summary of costs, including the minimum, average, median, and maximum values.
@@ 5 Evaluation
This section presents the gas costs evaluation of the proxy and diamond smart contract patterns, against a baseline built using the classic pattern in the context of an app deployment and upgrade  [^4] All results can be found here: https://anonymous.4open.science/r/proxy-diamond-patterns-gas-analysis/data.
(a) Deployment cost by pattern and version.
(b) Average function cost by pattern for version 2.
Figure 3: Evaluation results
5.1 Gas Cost During Smart Contract Deployment and Upgrades
@@
The primary aim is to compare the gas costs during the deployment (version 1) and upgrades (version 2 and 3) of the file notarization application implemented with the three different patterns. For the classic and proxy pattern, a new contract is deployed for each version. For the proxy, the proxy pointer is updated in the smart contract implementation. For the diamond pattern, in version 2, only facet is changed, and in version 3, all facets are changed.
Figure 3(a) presents the deployment costs, in gas units, of each version by pattern. The cost of deployment increases with each version, except for the diamond pattern, where the initial cost of deployment is significant. In all, the classic pattern requires just 1,614,545 units of gas, while the proxy and diamond patterns consume around 2.6 times as much, at 4,343,104 and 4,123,977 units of gas respectively. The classic pattern appears the most gas efficient. The increased consumption at deployment of the proxy and diamond patterns relates to the need to deploy more contracts compared to the classic pattern: two for the proxy pattern, and four for the diamond one.
5.2 Gas Cost During Smart Contract Execution
The analysis now shifts to comparing the average gas cost of functions to determine the most gas efficient execution pattern. Gas cost behaviors in versions 1 and 3 align with online gas reports. Notably, addFile and updateFile functions are the costliest, while remaining functions incur comparatively lower gas costs. Across the three patterns, the proxy and diamond patterns exhibit slightly higher expenses than the classic pattern. Notable cost discrepancies include compareHashes being significantly cheaper than addFile in the classic pattern. This difference is attributed to addFile requiring two storage writes while other functions mainly perform storage reads. The classic pattern avoids delegation to other contracts, unlike the proxy and diamond patterns, resulting in additional costs. Furthermore, incurring more operations when invoking a diamond adds to its expense compared to a proxy. Getters in the classic pattern involve state lookups, costing 2100 gas for cold storage loads, while compareHashes is a pure function, devoid of state lookups or modifications, resulting in lower gas costs.
@@ 5.2 Gas Cost During Smart Contract Execution
In reviewing the data obtained from the comparative analyses, the deployment phase exhibits the most significant variation among the different patterns, primarily due to the fluctuating number of smart contracts deployed. For this metric, the diamond pattern is the cheapest, especially for minor upgrades. During execution, the gas costs are similar across patterns, though there is a marginal escalation in costs when transitioning from the classic architecture to the proxy pattern, and subsequently from the proxy to the diamond pattern. This increase in gas costs is due to the additional operations required to delegate calls to the logic contract(s) in the proxy and diamond patterns.
6 Discussion
This section summarizes and consolidates the proxy and diamond patterns following the Alexandrian form format proposed by C. Alexander [2]. A decision model is proposed to help developers choose between both patterns.
6.1 Proxy Pattern Outline
@@
  • Summary: The proxy pattern facilitates upgradability in smart contracts through a proxy contract pointing to the latest version of a logic contract.
@@
  • Context: A smart contract must be upgraded due to evolving requirements and potential improvements [7].
@@
  • Problem: Traditional smart contracts lack the ability to be updated without manual storage migration, posing challenges in addressing vulnerabilities, enhancing functionality, and adapting to changing circumstances.
  • Forces (tradeoffs): The problem requires balancing the following forces: (i) Immutability vs. Upgradability. Smart contracts on blockchain platforms are traditionally immutable once deployed; (ii) Gas Costs vs. Flexibility. Minimizing gas costs while providing flexibility for contract upgrades; (iii) Trust vs. Transparency. Establishing trust in the upgrade process while maintaining transparency.
@@
  • Solution: The smart contract proxy pattern introduces a proxy contract as an intermediary layer. This proxy delegates calls to a logic contract, allowing seamless upgrades by deploying new logic contracts and updating the proxy to point to the latest version, hence allowing for dynamic updates.
  • Consequences:
@@
        Upgradability. Allows upgrades without the need to change the contract address nor requiring data migration;
        Simplest Upgradeable Pattern. The upgrade process consists in deploying a new logic contract and a proxy update pointing to it with a proxy administration function.
@@
        Compatibility Maintenance. Requires consistent function selectors and storage layouts;
        Limited Direct Function Visibility. Logic functions visibility only accessible in the documentation;
        Storage Collision. It requires careful consideration of storage layout to avoid storage overlap between the proxy and the logic contract as if both contracts use the same storage slot, it can lead to data loss. A convention is to use namespaced storage layouts for naming struct holding storage variables [18].
        Function Selector Clash. Different functions having the same selector can override each other [26]. This requires careful naming of function selectors as the Solidity compiler cannot detect function selector clashes between the proxy and logic contracts due to cross-contract interactions.
  • Related patterns: Diamond pattern.
  • Known uses: Two standardized implementations are proposed in the Ethereum Improvement Proposal EIP-897 and in OpenZeppelin’s smart contract development framework [22, 3]. Compound and USDC, respectively a DeFi protocol and a stable coin, both implement the proxy pattern for upgrades.
6.2 Diamond Pattern Outline
  • Summary: employing multiple implementation contracts to balance contract size, maintainability and versioning;
  • Context: Need for a solution to improve maintainability in large contracts.
  • Problem: Traditional approaches face challenges in managing contract size and versioning effectively.
  • Forces (tradeoffs): The problem requires balancing the following forces: (i) Contract Size vs. Modularity Balancing the need for compact contract sizes with the demand for modular, well-organized code structures; (ii) Maintainability vs. Simplicity Achieving improved maintainability without introducing unnecessary complexity; (iii) Scalability vs. Consistency Scaling smart contract applications requires accommodating versioning and updates.
  • Solution: The diamond pattern introduces a structure where a proxy can point to multiple logic contracts. It involves the deployment of all contracts (diamond and facets), retrieving facets function selectors, and leveraging both to implement a diamond cut.
  • Consequences:
@@ 6.2 Diamond Pattern Outline
        Better upgradability. Possibility to deploy smaller contracts during upgrades or updates to already deployed facets without requiring address change or data migration;
        Modularity. Code reusable across multiple contracts. A facet can be used in multiple diamonds;
        Contract size. Thanks to a modular structure, it can theoretically support an infinite number of facets. Therefore, the whole smart contract system has no size limit;
        Cheaper minor upgrades. Most of the time, only one facet is updated, so only small contracts are deployed for low gas costs;
        Shorter compilation time. Only modified facets need to be compiled, so for the same logic code, the compilation time is shorter than for the classic pattern and proxy pattern.
@@
        Implementation Complexity. A more complex structure compared to the classic and proxy patterns, and a lack of supporting libraries.
        Complexity in managing multiple logic contracts. Managing multiple logic contracts require careful consideration during upgrades. It requires developers to manage the diamond storage manually because of the multiple implementation contracts.
        Limited Direct Function Visibility. Like the proxy pattern, users depend on documentation to identify callable functions.
        Storage collision risks similarly to Proxy pattern
        Function selector clash similarly to Proxy pattern
  • Related patterns: Proxy pattern.
  • Known uses: Aavegotchi, a NFT-based gaming protocol; GeoWeb, a dApp managing digital land property rights using NFTs.
6.3 Choosing between the Proxy and Diamond Patterns
Figure 4: Decision model for upgradeable smart contract pattern usage.
Figure 4 proposes a decision model to orient between a proxy or a diamond pattern in the design stage of an upgradeable smart contract. It follows the design goal decision model introduced by Xu where decisions are modeled using a logical BPMN flow [32]. More precisely, arrows, logical gateways, and functional and non-functional properties orient the decision path.
The choice of smart contract pattern largely depends on the need for upgradability. For scenarios without upgradability needs, the classic pattern is preferable due to its straightforward development process, relying on contract inheritance and library imports. However, upgrades in the classic pattern often involve complex and resource-intensive data migration. This pattern also poses challenges in communicating new contract addresses to users, potentially affecting the user experience. For extensive upgradeable features, the diamond pattern is recommended. Its modular nature allows for easy addition or removal of facets, reducing compilation time. However, it is less cost-effective initially compared to the proxy pattern and requires in-depth knowledge of smart contract storage and facet-library management. The proxy pattern is advised for limited code sizes or infrequent upgrades. It simplifies development and integrates easily with libraries like OpenZeppelin’s. This pattern enhances upgradability by separating logic and state, reducing the need for data migration. But it offers less flexibility and modularity compared to the diamond pattern and demands careful consideration to maintain compatibility across versions.
In the end, while the classic pattern excels in execution ease, the proxy and diamond patterns provide a more manageable framework for upgrades, simplifying contract interactions for users during updates. The diamond pattern is more suitable for extensive upgrades, while the proxy pattern is recommended for limited upgrades and simpler development.
7 Conclusion
@@
Each pattern unfolds with distinct strengths and weaknesses, delineating its applicability across diverse scenarios. The classic pattern implies an unpractical and costly approach to smart contract upgrades because of the need of manual data migration. The proxy pattern offers the simplest solution for upgradability, but security concerns such as storage collisions and function selector clashes remain. This demands a developer’s careful attention and thus results in a more challenging pattern to utilize. The diamond pattern is the most complex of the three, but it offers the most flexibility and maintainability thanks to its modularity. Moreover, the diamond pattern is the most gas efficient when it comes to doing more than two upgrades. This is because the diamond pattern does not need to deploy long contracts, but only small facets. Finally, this pattern is the most scalable because it does not have a contract size limit. Despite these considerations, real-world implementations in projects like Compound, USDC, GeoWeb, and Aavegotchi underline the use of the proxy and diamond patterns as facilitators for flexible, upgradeable, and scalable smart contracts.
To generalize these initial findings, the study advocates for extending experiments beyond the notarization scenario used as a comparison baseline for this paper. Essential to this endeavor is the necessity for additional experiments encompassing diverse blockchain networks and pattern libraries to extrapolate and validate the findings. For future work, a replication of the study on alternative upgradeable patterns would also provide a more comprehensive understanding of upgradeable smart contract patterns.